a force 2N is applied from a point O to perpendicular distance 4m.find the torque due to this force.
1
Expert's answer
2020-09-21T08:30:59-0400

SOLUTION

given data

force applied(F)=2N

perpendicular distance(d)=4m

torque can be given as

"\tau=Fdsin\theta"

force is perpendicular to distance so angle wiil be "\frac{\pi}{2}rad"


"\tau=2\times4sin\frac{\pi}{2}"


"\tau=8Nm"